derive equation of motion for uniformly accelerated .what will be its speed at the end of 100s?
1
Expert's answer
2012-05-22T08:22:31-0400
s=s0+v0t+at2/2 whete s0-initial displacement, v0-initial velosity, a=const - acceleration
if we derive this equation we'll get
v=v0+at
So if we need to know the speed when t=100s we substitute t=100 into the last equation and obtain
v=v0+100a